OpenStudy (anonymous):

Describe how the graph of y = x2 can be transformed to the graph of the given equation. y = (x-16)2 Shift the graph of y = x2 down 16 units. Shift the graph of y = x2 left 16 units. Shift the graph of y= x2 right 16 units. Shift the graph of y = x2 up 16 units.

OpenStudy (anonymous):

when you have y = (x - k)^2, it means you have y=x^2 shifted to the right k units

OpenStudy (anonymous):

to confirm this, put in some values for x and see how they compare to a normal x^2 parabola